1. Let f be the function defined by

f (x) = x2

-------------

(2x + 1)2

:

(a) Determine the vertical and horizontal asymptotes (show all limits). (4)

(b) Use the sign pattern for f ' (x) to determine

(i) the interval(s) over which f rises and where it falls; (4)

(ii) the local extrema. (2)

(c) Use the sign pattern for f " (x) to determine

(i) where the graph of f is concave up and where it is concave down. (4)

(ii) the inflection points (if any) (2)


1
Expert's answer
2022-04-22T14:00:44-0400

"f(x)=\frac{x^2}{(2x+1)^2}" , "x\neq -\frac{1}2"


(a) Vertical asymptotes: "(2x+1)^2=0,\quad 2x+1=0,\quad x=-\frac{1}{2}" .

"\lim\limits_{x\rightarrow -1/2}\frac{x^2}{(2x+1)^2}=\infty ,\quad" "x=-\frac{1}{2}" is a vertical asymptote.


Horizontal asymptotes: "\lim\limits_{x\rightarrow\infty}\frac{x^2}{(2x+1)^2}= \lim\limits_{x\rightarrow\infty}\frac{x^2}{4x^2+4x+1}= \lim\limits_{x\rightarrow\infty}\frac{1}{4+4/x+1/x^2}=\frac{1}{4}"

"y=\frac{1}{4}" is a horizontal asymptote.

(b) "f^\prime (x)=\frac{2x\cdot (2x+1)^2-4(2x+1)\cdot x^2}{(2x+1)^2}=\frac{2x}{(2x+1)^3}"


(i) If "-\tfrac{1}{2}<x<0" , then "f^\prime (x)<0" and "f(x)" falls.

If "x<-\frac{1}{2}" or "x\geq 0" , then "f^\prime (x)>0" and "f(x)" rises.


(ii) "f(x)=0" , "\frac{2x}{(2x+1)^3}=0" , "x=0" .

The point "x=0" is a local minimum (if "x<0" , then "f" falls; if "x>0" , then "f" rises).


(c) "f^{\prime \prime}(x)=\frac{2\cdot (2x+1)^3-6(2x+1)^2\cdot 2x}{(2x+1)^6}=\frac{2-8x}{(2x+1)^4}"


(i) If "x<\frac{1}{4}" , then "f^{\prime\prime}(x)>0" and "f(x)" is concave up.

If "x>\frac{1}{4}" , then "f^{\prime\prime}(x)<0" and "f(x)" is concave down.


(ii) "f^{\prime\prime}(x)=0,\ \ \frac{2-8x}{(2x+1)^4}=0,\ \ x=\frac{1}{4}" .

The point "x=\frac{1}{4}" is an inflection point (if "x<\frac{1}{4}" , then "f" is concave up; if "x>\frac{1}{4}" , then "f" is concave down).